# DeepVideo-R1: Video Reinforcement Fine-Tuning via Difficulty-aware Regressive GRPO

### v3

@@ 2 Related works
Large Language Models (LLMs) [brown2020language, yang2024qwen2, grattafiori2024llama] have exhibited strong generalization and reasoning capabilities across a wide range of domains, including knowledge-intensive tasks [jin2025search, luo2025graph, park2023graph, park2024generative], mathematical reasoning [shao2024deepseekmath, luo2023wizardmath], and scientific domain [zheng2025large, bai2025intern, park2024llamo]. With the strong reasoning abilities of LLMs, Video Large Language Models (Video LLMs) have sextend the reasoning capabilities of LLMs to dynamic video domains, achieving notable performance on various tasks [wu2024longvideobench, fu2024video, li2024mvbench, shen2024longvu], such as video question-answering [bai2025qwen2, zhu2025internvl3, zhang2025videollama, li2023videochat, ko2023large, li2024llava] and video captioning [xu2023mplug, cheng2024videollama, zhang2024video], through a comprehensive understanding of video content. Despite their impressive performance, Video LLMs exhibit limited performance on long video inputs or fine-grained video understanding tasks which require detailed spatial-temporal reasoning within videos [gao2017tall, krishna2017dense, xiao2024can]. Most existing methods primarily emphasize video perception or short-context understanding, often relying on static supervised fine-tuning objectives that fail to capture reasoning dynamics. To address these challenges, we leverage a reinforcement learning-based fine-tuning approach to improve the reasoning and generalization capabilities of VideoLLMs

### v4

@@ 2 Related works
Large Language Models (LLMs) [brown2020language, yang2024qwen2, grattafiori2024llama] have exhibited strong generalization and reasoning capabilities across a wide range of domains, including knowledge-intensive tasks [jin2025search, luo2025graph, park2023graph, park2024generative], mathematical reasoning [shao2024deepseekmath, luo2023wizardmath], and scientific domain [zheng2025large, bai2025intern, park2024llamo]. With the strong reasoning abilities of LLMs, Video Large Language Models (Video LLMs) have extended the reasoning capabilities of LLMs to dynamic video domains, achieving notable performance in various tasks [wu2024longvideobench, fu2024video, li2024mvbench, shen2024longvu], such as video question-answering [bai2025qwen2, zhu2025internvl3, zhang2025videollama, li2023videochat, ko2023large, li2024llava] and video captioning [xu2023mplug, cheng2024videollama, zhang2024video], through a comprehensive understanding of video content. Despite their impressive performance, Video LLMs exhibit limited performance on long video inputs or fine-grained video understanding tasks which require detailed spatial-temporal reasoning within videos [gao2017tall, krishna2017dense, xiao2024can]. Most existing methods primarily emphasize video perception or short-context understanding, often relying on static supervised fine-tuning objectives that fail to capture reasoning dynamics. To address these challenges, we leverage a reinforcement learning-based fine-tuning approach to improve the reasoning and generalization capabilities of VideoLLMs